OpenStudy (ulahlynn):

A line passes through (1,9) and is parallel to the line through (-2,2) and (4,5). Find its equation.(general form). PLEASE HELP!

OpenStudy (anonymous):

First you have to find the slope of a line passing through points (-2,2) and (4,5). Formula for slope is: \(\Large m=\frac{y_2-y_1}{x_2-x_1}\) Since the line you are looking for is parallel to the line of these points, it will have the same slope with the slope of the these two points. Next thing you have to do is use the point (1,9) and the slope pulg these values in this equation: \((y-y_1)=m(x-x_2)\). Then simplify. Remember: The general form of a line is \(Ax+By+C=0\). So you have to rearrange the equation you got to form the general equation of a line.

OpenStudy (ulahlynn):

ohh. thank you soo much AntiNode :) can you pls. help me with this last question. Find The equation of the perpendicular Bisector of a line joining (4,0) and (-6,-3)?

OpenStudy (anonymous):

Biseector is a line that divides a line into two equal parts. First, find the midpoint of the two points: \(midpoints=\Large (\frac{x_2+x_1}{2}, \frac{y_2+y_1}{2})\) So the line that you are looking for will pass through this point. Next thing you have to do is find the slope of the two given points. The \(\color{blue}{negative\ reciprocal}\) of this slope will be the slope of the required line. Since you already have a point and a slope, you can easily apply the same method that you did on the first question.

OpenStudy (ulahlynn):

oohhhh i see... thank youu soo much again AntiNode!!.. :D

OpenStudy (anonymous):

You're welcome.